### v2

# distributive quandle and detecting causality

Abstract
I investigated the capability of a specific type of quandle, quandle whose operation satisfying that $(a_{1}*b_{1})*(a_{2}*b_{2})=(a_{1}*a_{2})*(b_{1}*b_{2})$, to detect causality in (2+1)-dimensional globally hyperbolic spacetime by determining if they can distinguished the connected sum of two Hopf links with an infinite series of relevant three-component links constructed by Allen and Swenberg in 2020, who suggested that any link invariant must be able to distinguish those links for them to detect causality in the given setting. I show that these quandles fail to do so as long as $a\sim b\Leftrightarrow a*b=a$ defines an equivalence relation. The Alexander quandles is an example that this result can apply to.
1 Introduction
Mathematical knots are embeddings of circles into 3-dimension Euclidean space. Two knots are equivalent only if they can be transformed to each other via a series of Reidemeister moves. A Link is a collection of knots that are intertwined together. A knot or link invariant is a function that assign same value to equivalent knots/links, which require it to be preserved under all Reidemeister moves. Using link invariant to distinguish link has been shown to relates detection of causality in spacetime with the following work from former researchers:
Let $X$ be a $(2+1)$-dimensional globally hyperbolic spacetime with Cauchy surface $\sum$ homomorphic to $\mathbb{R}^{2}$. Chernov and Nemirovski[1] proved the Low conjecture[2], which state that as long as $\sum$ is not homomorphic to $S^{2}$ or $\mathbb{R}p^{2}$, then two events $x,y\in X$ are causally related if and only if their skies $S_{x}\cup S_{y}$ are topologically linked. Two causally unrelated points in spacetime give a link isotopic to the connected sum of two Hopf link, denoted as $L_{2H}$ in this paper. This result gave rise to the possibility that the causality of two events in their skies can be detected by link invariants.
Allen and Swenberg[3] investigate the natural question that whether $S_{x}$ and $S_{y}$, and thus the causality, can be detected by link invariant. If there exist a three component link consists of one unknot and other two individually deformable to the longitudes of the solid torus $S^{1}\times R^{2}$ whose link invariant is the same as $L_{2H}$, then it is said that the link invariant can not completely determine causality in $X$. They showed that Conway-polynomial cannot detect causality in that space time, as there exist a infinite family of three-components two-sky-link link (Allen-Swenberg links) that can not be distinguished with the connected sum of two Hopf links by Conway polynomial.
Section 2 of this paper introduced link invariant generated by quandle, an algebraic structure, and specifically for a specific type of quandle whose operation satisfy the property $(a_{1}*b_{1})*(a_{2}*b_{2})=(a_{1}*a_{2})*(b_{1}*b_{2})$ for all $a_{1},b-1,a_{2},b_{2}$ in the set, which I call the distributive quandle. In section 3-4, this paper investigate the capability of distributive quandles for distinguishing $L_{2H}$ and Allen-Swenberg links. I show that as long as the relation define on a distributive quandle by $a~{}b\Leftrightarrow a*b=a$ is an equivalence relation, then it cannot distinguish $L_{2H}$ and any of Allen-Swenberg link in the series, thus fail to capture causality in the spacetime. Since Alexander quandles satisfy the restriction on qunadle in this result, an application of this result is that all Alexander quandle fail to detect causality in the given spacetime.
2 Mathematical background
definition 2.1: quandle
A quandle $Q(*)$ a set $Q$ equipped with a binary operation $*:Q^{2}\to Q~{}|~{}(x,y)\to x*y$ such that the operation satisfy the following requirement: $\forall x,y,z\in Q:$
1. $x*x=x$
2. There exist an unique $q\in Q$ such that $q*y=x$. We can then define the inverse operation $*^{-1}$ as $x*^{-1}y=q$.
3. $(x*y)*z=(x*z)*(y*z)$
definition 2.2: quandle homomorphism
Given two quandle $A(*)$ and $B(\times)$, a homomorphism from $A(*)$ to $B(\times)$, $f\in Hom(A(*),B(\times))$, is a function $f:A\to B$ such that $f(x*y)=f(x)\times f(y)~{}\forall x,y\in A$.
definition 2.3 fundamental quandle of a link
For an oriented knot or link diagram with arc-set $L$, its fundamental quandle $Q(L,*)$ is the quandle on $L$ with quandle operation defined by crossing relations given by each crossing in the link diagram in the following way:
Theorem 2.1
The fundamental quandle is a link variant.
Miller[4] proved this by showing that quandle axioms are motivated by the Reide- meister moves in such a way that the fundamental quandle is locally invariant.
definition 2.4 quandle coloring invariant
For a link $L$ with its fundamental quanlde $Q(L,*)$ and a finite quandle $X(\times)$, the coloring space is set of all homomorphisms from $Q(L,*)$ to $X(\times)$, $Hom(Q(L,*),X(\times))$. We say an element in the space is a coloring on the link by quandle $X$. It can be regarded as a way of assigning an element in $X$ to each arc in $L$ such that those elements satisfy the crossing relations defined in $L$ but with the quandle operation in $X(\times)$.
The quandle coloring invariant is $|Hom(Q(L,*),X(\times))|$, the cardinality of coloring space.
Remark: The coloring space from a finite quandle $X$ on any links or knots $Q(L)$ always contains the set $C_{T}=\{f:Q(L)\to X,~{}a\to x~{}|~{}x\in X\}$, namely, the set of constant function from $Q(L)$ to $x$, where $|C_{T}|=|X|$. An element in $C_{T}$ is called the trivial coloring, sending every arcs in the link or knot to the same element in the quandle.
definition 2.5: enhanced coloring invariant
Defined by Nelson[6], using the same setting in definition 2.4, the enhanced coloring invariant is a polynomial in $q$ assign to the coloring space:
$\Phi(L,X)=\sum_{f\in Hom(Q(L),X(\times)}q^{|IM(f)|}$
This enhanced link invariant contains strictly more information then the quandle coloring invariant defined in 2.4.
definition 2.6 distributive quandle
Consider the type of quandle $Q(*)$ with the property that $\forall x_{1},y_{2},x_{2},y_{2}\in Q$, they satisfy that $(x_{1}*y_{2})*(x_{2}*y_{2})=(x_{1}*x_{2})*(y_{1}*y_{2})$. Let’s call this type of quandle the distributive quandle. Let $*^{-1}$ denote its inverse operation. It can be proved that a distributive quandle satisfies the following 9 properties, for all $x,y,z,x_{1},x_{2},y_{1},y_{2}\in Q$:
$x*(y*z)=(x*x)*(y*z)=(x*y)*(x*z)~{}(1)$
$[(x_{1}*^{-1}y_{1})*(x_{2}*^{-1}y_{2})]*(y_{1}*y_{2})=[(x_{1}*^{-1}y_{1})*y_{1}]*[(x_{2}*^{-1}y_{2}*y_{2})]=x_{1}*x_{2}\to(x_{1}*^{-1}y_{1})*(x_{2}*^{-1}y_{2})=(x_{1}*x_{2})*^{-1}(y_{1}*y_{2})~{}(2)$
$(2)$ implies $x*^{-1}(y*z)=(x*x)*^{-1}(y*z)=(x*^{-1}y)*(x*^{-1}z)~{}(3)$, and
$[(x_{1}*^{-1}y_{1})*^{-1}(x_{2}*^{-1}y_{2})]*(y_{1}*^{-1}y_{2})=_{(2)}x_{1}*^{-1}x_{2}\to$
$(x_{1}*^{-1}y_{1})*^{-1}(x_{2}*^{-1}y_{2})=(x_{1}*^{-1}x_{2})*^{-1}(y_{1}*^{-1}y_{2})~{}(4)$
$(x*^{-1}y)*^{-1}z=(x*^{-1}y)*^{-1}(z*^{-1}z)=(x*^{-1}z)*^{-1}(y*^{-1}z)~{}(5)$ $x*^{-1}(y*^{-1}z)=(x*^{-1}x)(y*^{-1}z)=(x*^{-1}y)*^{-1}(x*^{-1}z)~{}(6)$
$(x*^{-1}y)*z=(x*^{-1}y)*(z*^{-1}z)=(x*z)*^{-1}(y*z)~{}(7)$
$x*(y*^{-1}z)=(x*^{-1}x)*(y*^{-1}z)=(x*y)*^{-1}(x*z)~{}(8)$
$[(x*^{-1}z)*(y*^{-1}z)]*z=x*y\to$ $(x*y)*^{-1}z=(x*^{-1}z)*(y*^{-1}z)~{}(9)$$(9)$ is true for all quandle)
Example 2.7: Alexander quandle
An Alexander quandle is a module $M$ over $\mathbb{Z}[t^{\pm}]$ under the the operation $a*b=ta+(1-t)b$. Since we need finite quandle for quandle coloring link invariant, we consider finite Alexnader quandle which is in the form $Q=\mathbb{Z}_{n}[t^{\pm}]/(h(t))$ where $h$ is a monic polynomial in $t$ (Nelson[5]). It is a distributive quandle because $\forall a_{1},b_{1},a_{2},b_{2}\in Q$,
$(a_{1}*b_{1})*(a_{2}*b_{2})=(ta_{1}+(1-t)b_{1})*(ta_{2}+(1-t)b_{2})=(t^{2}a_{1}+t(1-t)b_{1})+(t(1-t)a_{2}+(1-t)^{2}b_{2})=(t^{2}a_{1}+t(1-t)a_{2})+(t(1-t)b_{1}+(1-t)^{2}b_{2})=(ta_{1}+(1-t)a_{2})*(tb_{1}+(1-t)b_{2})=(a_{1}*a_{2})*(b_{1}*b_{2})$.
2.1 Example: coloring on the trefoil knot
The trefoil knot (denoted as $k_{3}$) is the following:
The fundamental quandle of $k_{3}$ is in the form $Q(k_{3},\times)=\{a,b,c~{}:~{}c\times b=a,b\times a=c,a\times c=b\}$. Consider the quandle on set $\mathbb{Z}_{3}$ with operation $x*y=2x-y$. Then to derive the coloring space of this quandle on the trefoil knot is essentially solving this system of modulo equations: $\{2c-b=a,2b-a=c,2a-c=b~{}(mod~{}3)\}$. One can check that the solutions require that $a=b=c,a\in\mathbb{Z}_{3}$.
This means $Hom(Q(k_{3},\times),\mathbb{Z}_{3}(*))=\{(a\to 1,b\to 1,c\to 1),(a\to 2,b\to 2,c\to 2),(a\to 3,b\to 3,c\to 3)\}$. The coloring invariant from $\mathbb{Z}_{3}(*)$ on the trefoil knot is therefore $|~{}Hom(Q(k_{3},\times),\mathbb{Z}_{3}(*))~{}|=3$. All function has only 1 image, so the enhanced polynomial is $\Phi(k_{3},\mathbb{Z}_{3}(*))=3q^{1}$.
3 distributive quandle & detecting causality
Allen and Swenberg[2] investigated whether there are link invariants can detect causality in $X$, a (2+1)-dimensional globally hyperbolic spacetime with a Cauchy surface $\sigma$ whose universal cover is homomorphic to $R^{2}$. They constructed an infinite series of three-component links (Allen-Swenberg links), and suggested that such link invariant needs to be able to distinguish between these links and connected sum of two Hopf ($L_{2H})$. (see figures of the links below). They found that Conway polynomial fail to do so. I would show the coloring invariant and its enhanced polynomial from any distributive quandle with $a\sim b\Leftrightarrow a*b=a$ defining an equivalence relation also can’t differentiate $L_{2H}$ from the Allen-Swenberg links.
Figure 3.1: connected sum of two Hopf links,denoted by $L_{2H}$
The first Allen-Swenberg link in series, $A_{1}$ is
The nth Allen-Swenberg link, denoted by $A_{n}$, is form by duplicating the complex tangle in the middle component :
Figure 3.2: general Allen-Swenberg links $A_{n}$ and the tangle $T$ that consists of it
3.1 Proposition 3.1
Let $Q(*)$ be a distributive quandle (definition 2.6) such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*b=a$ is an equivalence equation.
Let $C_{A_{i}}$ and $C_{H}$ denote the cardinality of the coloring space of the quandle on the ith Allen-Swenberg link and the connected sum of two Hopf links, respectively (i.e., $C_{A_{i}}=|Hom(A_{i},Q(*))|,~{}C_{H}=|Hom(L_{2H},Q(*))|$. Let $\Phi(A_{i},Q(*),\Phi(L_{2H},Q(*))$ denote the enhanced coloring polynomial(definition 2.5) of the two links under the coloring by $Q(*)$, respectively. Then
$C_{A_{i}}=C_{H}$ and $\Phi(A_{i},Q(*))=\Phi(L_{2H},Q(*)),~{}\forall i\in\mathbb{Z}^{+}$.
This means any distributive quandle with the equivalence relation defined above is unable to distinguish between the any of Allen-Swenberg links and the connected sum of Hopf link by their coloring invariant or the enhanced version, therefore fail to detect causality in the given setting.
Example: Alexnader quandle
All finite Alexander quandle defined in section 2.7 satisfy the property of $Q(*)$ defined above. Section 2.7 show it is distributive. The relation $\sim$ defined on a finite Alexander quandle $A=\mathbb{Z}_{n}[t^{\pm}]/(h(t))$ is
$a\sim b\Leftrightarrow a*b=a\Leftrightarrow ta+(1-t)b=a~{}(mod~{}h(t))$. This is equivalent to
$(t-1)a=(t-1)b~{}(mod~{}h(t))$, which is clearly an equivalence relation. So proposition 3.1 imply that Alexnader quandle are unable to detect causality in the given spacetime.
4 Proof for proposition 3.1
First we prove some properties of distributive quandle which will be used later. In some calculation process below, $=_{(i)}$ or $=_{k}$ means that the right side of the equation are derived by applying property $(i)$ in section 2.6 or property k in this section that is already proved. Let $Q(*)$ be a distributive quandle. Then
1.$[(a*x)*^{-1}y]*z=[(a*z)*^{-1}y]*x,~{}\forall a,x,y,z\in Q$.
Proof :
$[(a*x)*^{-1}y]*z=_{(7)}[(a*x)*z]*^{-1}(y*z)=[(a*z)*(x*z)]*^{-1}(y*z)=_{(2)}[(a*z)*^{-1}y]*[(x*z)*^{-1}z]=[(a*z)*^{-1}y]*x$, as desired.
2. $[(a*^{-1}x)*y]*^{-1}z=[(a*^{-1}z)*y]*^{-1}x$.
Proof: A easy way to proof this is noticing that the orignal and inverse operation * and $*^{-1}$ is completely symmetric in a distributive quandle. This is because one can check that replacing all * by $*^{-1}$ and all $*^{-1}$ by * in any one of properties in section 2.6 gives another property that is alsp satisfied by $Q(*)$. So if an equation for operation in $Q(*)$ holds, than the equation derived by switching between all * and $*^{-1}$ in the orignal equation is also valid. By doing it to 1 we get $[(a*^{-1}x)*y]*^{-1}z=[(a*^{-1}z)*y]*^{-1}x$, which is exactly 2, which complet the proof.
definition 4.1
For any $x,y\in Q$, define $f_{xy}:Q\to Q$ by $f_{xy}(q)=(q*^{-1}x)*y$.
Let $f_{x_{2}y_{2}}\circ f_{x_{1}y_{1}}(q):=f_{x_{2}y_{2}}(f_{x_{1}y_{1}}(q))$ denote the composition of two such function. Then $\forall x,y,a,b,q\in Q$,
4. $f_{xy}\circ f_{ab}=f_{ab}\circ f_{xy}$. Proof: $f_{xy}\circ f_{ab}(q)=[[(q*^{-1}a)*b]*^{-1}x]*y=_{1}[[(q*^{-1}x)*b]*^{-1}a]*y=_{2}[[(q*^{-1}x)*y]*^{-1}a]*b=f_{ab}\circ f_{xy}(q)$
5. $f_{1}\circ f_{2}\circ...\circ f_{k}=f_{1^{\prime}}\circ f_{2^{\prime}}\circ...\circ f_{k^{\prime}}$, where $\{i\}\to\{i^{\prime}\}$ is a permutation, and $f_{i}=f_{x_{i}y_{i}}$ for some $x_{i},y_{i}\in Q$.
Proof: every permutation can by derived by a series of transposition, and each transposition can be derived by a series of switch between adjacent element. It follows from 4 that switching between adjacent individual functions preserve the composited function, which complete this proof.
6. $f_{xy}(a*b)=f_{xy}(a)*f_{xy}(b)$ and $f_{xy}(a*^{-1}b)=f_{xy}(a)*^{-1}f_{xy}(b)$. Proof:
$f_{xy}(a*b)=[(a*b)*^{-1}x]*y=[(a^{-1}x)*(b*^{-1}x)]*y=[(a^{-1}x)*y]*[(b^{-1}x)*y]=f_{xy}(a)*f_{xy}(b)$, and
$f_{xy}(a*^{-1}b)=[(a*^{-1}b)*^{-1}x]*y=[(a*^{-1}x)*^{-1}(b*^{-1}x)]*y=[(a*^{-1}x)*y]*^{-1}[(b^{-1}x)*y]=f_{xy}(a)*^{-1}f_{xy}(b)$.
4.1 coloring on connected sum of two Hopf links
Let $Q(*)$ be a distributive quandle such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*c=a$ is an equivalence equation. For the connected sum of two Hopf links, the system of equations associated to the coloring space comes from only 4 crossings, so it can be calculated relatively easily. Consider the following label by a valid coloring on $L_{2H}$ by $Q$:
The four crossing relation give the following requirements:
$t_{1}:$ $x*y_{1}=x$. Since $\sim$ is an equivalence equation, this imply $x\sim y_{1}\to y_{1}\sim x\to y_{1}*x=y_{1}$
$t_{2}:$ $y_{2}=y_{1}*x=y_{1}.$. So $y_{1}=y_{2}$
$t_{3}:$ $y_{1}=y_{2}*z\to y_{1}*z=y_{1}$, so $y_{1}\sim z$
The first three crossings generate the solution set $\{x,y_{1},y_{2},z\}=\{x,y,y,z\}$ with the condition that $x\sim y\sim z$ where $x,y,z\in Q$. It automatically satisfy the equation given by $t_{4}:$ $z*y_{2}=z$ because we have $y_{1}=y_{2}$ and $y_{1}\sim z$ which imply $z\sim y_{2}\to z*y_{2}=z$.
4.2 Lemma 1
Consider the following tangle as a subpart of the Allen-Swenberg link (called $T$ in section 3) , which I will call $L_{p}$.
Claim: when applying coloring from any distributive quandle $Q(*)$, the system of equations with the operation from $Q(*)$ given by this subpart link diagram force $n=d$ and $m=e$.
4.3 proof of lemma 1
Firstly, notice that the following subpart of link diagram $L_{p^{\prime}}$ is mathematically the same as $L_{p}$ because they can transform to each other by only Reidemeister move type $III$:
Figure 4.1: Reidemeister moves type $III$
transformation:
This mean it suffices to prove the the two pairs of exiting line segment of $L_{p}^{\prime}$ has the same value respectively.
Let $Q(*)$ be an arbitrary distributive quandle. Let’s first analysis the coloring of $Q(*)$ on the following two structures, called structure ① and ②,respectively. In both structures by labeling 4 of the arcs exposed outside with $x,y,n,m$ representing some element in $Q(*)$, the other 4 out-exposed aces labeled by $x^{\prime},y^{\prime},n^{\prime},m^{\prime}$ can be expressed by a string of quandle operations between $x,y,n,m$. By using the properties of distributive quandle given in section 2.6, these expressions can be further simplified to the form of some functions $f_{pq},~{}p,q\in\{x,y,n,m\}$ defined in definition 4.1 acting on one of $x,y,n,m$ The result is shown in the figure below and the calculation process is shown below.
Figure 4.2: structure 1(left) and structure 2(right). Recall the definition of $f_{xy}$ in definition 4.1
For structure ①:
$y^{\prime}=y*x=(y*^{-1}y)*x=f_{yx}(y)$
$x^{\prime}=x*^{-1}y^{\prime}=x*^{-1}(y*x)=(x*^{-1}y)*(x*^{-1}x)=(x*^{-1}y)*x=f_{yx}(x)$
$n^{\prime}=i*^{-1}x=(n*y^{\prime})*^{-1}x=[n*(y*x)]*^{-1}x=[(n*^{-1}x)*[(y*x)*^{-1}x]]=(n*^{-1}x)*y=f_{xy}(n)$.
$m^{\prime}=j*y^{\prime}=(m*^{-1}x^{\prime})*(y^{\prime})=[m*^{-1}[(x*^{-1}y)*x]]*(y^{\prime})=(m*y^{\prime})*^{-1}[[(x*^{-1}y)*x]*y^{\prime}]=(m*y^{\prime})*^{-1}[[(x*^{-1}y)*x]*(y*x)]=[m*(y*x)]*^{-1}x=(m*^{-1}x)*[(y*x)*^{-1}x]=(m*^{-1}x)*y=f_{xy}(m)$.
For structure ②: $x_{1}=x*y~{}~{}x_{2}=x_{1}*^{-1}m~{}~{}y_{1}=y*^{-1}m~{}~{}y_{2}=y_{1}*^{-1}x_{2}$, so
$x^{\prime}=x_{2}*n=(x_{1}*^{-1}m)*n=[(x*y)*^{-1}m]*n=[[(x*^{-1}x)*y]*^{-1}m]*n=f_{mn}\circ f_{xy}(x)$
$y^{\prime}=y_{2}*n=(y_{1}*^{-1}x_{2})*n=[(y*^{-1}m)*^{-1}(x_{1}*^{-1}m)]*n=[(y*^{-1}x_{1})*^{-1}m]*n=[[y*^{-1}(x*y)]*^{-1}m]*n=[[(y*^{-1}x)*y]*^{-1}m]*n=f_{mn}\circ f_{xy}(y)$
Now we can view $L_{P^{\prime}}$ as the combination of two structures 1 and one structure 2:
By applying the calculation result in Figure 4.2 on the three structures 1 or 2 in above figure one by one, we have
$d=f_{ji}(j),~{}j=f_{m^{\prime}n^{\prime}}\circ f_{ba}(b),~{}b=f_{yx}(y),~{}y=f_{ij}(n^{\prime}),~{}n^{\prime}=f_{xy}(n)$. and
$e=f_{ji}(i),~{}i=f_{m^{\prime}n^{\prime}}\circ f_{ba}(a),~{}a=f_{yx}(x),~{}x=f_{ij}(m^{\prime}),~{}m^{\prime}=f_{xy}(m)$. In order to simplify the expression for $j$, notice that these equations gives $b=f_{yx}\circ f_{ij}(n^{\prime}),~{}a=f_{yx}\circ f_{ij}(m^{\prime})$. Let $g:=f_{yx}\circ f_{ij}$. By property 5, $g((p*q))=g(p)*g(q),~{}g((p*^{-1}q))=g(p)*^{-1}(q)~{}~{}\forall p,q\in Q$. So $j=f_{m^{\prime}n^{\prime}}\circ f_{ba}(b)=f_{m^{\prime}n^{\prime}}[(g(n^{\prime})*^{-1}g(n^{\prime}))*g(m^{\prime})]=f_{m^{\prime}n^{\prime}}(g(n^{\prime}*m^{\prime}))=g(f_{m^{\prime}n^{\prime}}(n^{\prime}*m^{\prime}))=g([(n^{\prime}*m^{\prime})*^{-1}m^{\prime}]*n)=g(n^{\prime})=b$.
Similarly since $i=f_{m^{\prime}n^{\prime}}\circ f_{ba}(a)$, using similar method we can get $i=a$.
Pluging in $j=b,~{}i=a$, the above equations gives $d=f_{ji}\circ f_{yx}\circ f_{ij}\circ f_{xy}(n)$ and $e=f_{ji}\circ f_{yx}\circ f_{ij}\circ f_{xy}(m)$.
Note that $f_{vw}\circ f_{vw}$ is the identity function for all $v,w\in Q(*)$ because $\forall q\in Q(*),$
$f_{vw}\circ f_{vw}(q)=[[(q*^{-1}v)*w]*^{-1}w]*v=(q*^{-1}v)*v=q$. So combining these equations and rearranging each individual functions which property 5 (see at the start of section 4) allows, we have
$d=f_{ji}\circ f_{ij}\circ f_{yx}\circ f_{xy}(n)=f_{ji}\circ f_{ij}(n)=n$ and $e=f_{ji}\circ f_{ij}\circ f_{yx}\circ f_{xy}(m)=f_{ji}\circ f_{ij}(m)=m$.
This show under a valid coloring from any distributive quandle,the labeled arcs must satisfy the relation $d=n,e=m$, which complete the proof.
4.4 coloring on the first Allen-Swenberg link
Proposition 4.4
Let $Q(*)$ be a distributive quandle such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*c=a$ is an equivalence equation. Then referring to figure 4.3, its coloring space on the first Allen-Swenberg link $A_{1}$ is the solution set $\{x1=x2=x3=x4=c_{1},y1=y2=c_{2},~{}others=c_{3}\}$, where $c_{1},c_{2},c_{3}\in Q$ with the condition that $c_{1}\sim c_{2}\sim c_{3}$$others"$ refer to all line segment in figure 3 other than $x1,x2,x2,x4,y1,y2$.
Figure 4.3: $A_{1}$ with orientation. Some arc and crossing are labeled for writing the proof
Proof for proposition 4.4
Lemma 1 requires that $d_{1}=r,e_{1}=d_{2}=m,e_{2}=k.$ The crossing relations given by the 5 crossings on the top in Figure 4.3 gives:
$t_{2}\&t_{3}:$ $r=m*^{-1}x_{1}=k$. $t_{1}\&t_{4}:$ $x_{2}=x_{1}*r=x_{1}*k=x_{3}$. $t_{5}:$ $x_{4}=x_{2}*x_{3}=x_{3}*x_{3}=x_{3}$. So we can let $x:=x_{4}=x_{2}*x_{3}=x_{3}*x_{3}=x_{3}$. The 3 crossings at the bottom give:
$t_{6}:$ $x*y_{1}=x\to x\sim y_{1}\to y_{1}\sim x_{1}.$
$t_{7}:$ $y_{2}=y_{1}*^{-1}x_{1}$. Since $y_{1}\sim x_{1},~{}y_{1}*x=y_{1}\to y_{1}=x*^{-1}y_{1}$, so $y_{2}=y_{1}$. .
$t_{8}:$ $y_{1}*r=y_{2}\to y_{1}*r=y_{1}\to y_{1}\sim r$. Getting back to $t_{1}:$, we have $x_{1}=x*^{-1}r.$ Since by calculation above we have $y_{1}\sim r,~{}x\sim y_{1}$, so $x\sim r\to x*r=x\to x=x*^{-1}r$. So $x_{1}=x*^{-1}r=x$.
By $t_{2}:$, Since $x\sim r$ imply $r\sim x$, we have $m=r*x_{1}=r*x=r$.
Since $r=k,r=m$, we have $m=k=r$.
Since the modulo equations generated by existing crossing already required $r=m=k$, the complex middle parts of strands degenerated and became equivalent as the unknot:
This means that all variable in the middle part has to have the same value, which is $m$. Let $y:=y_{1},~{}z:=m$. Combining the results derived above that $x_{1}=x_{2}=x_{3}=x_{4},y_{1}=y_{2}$, $x\sim y_{1}\sim r=m$, the solution set for coloring from $Q(*)$ to $A_{1}$ should be $\{x1=x2=x3=x4=x,y1=y2=y,~{}others(the~{}middle~{}2~{}L_{p}~{}structures)=z\}$ with condition that $x\sim y\sim z$.
4.5 coloring on general Allen-Swenberg links
In the oriented $A_{k}$ digram, we can apply lemma 1 to label all arcs in the middle component outside each $T_{i}$ by $m_{j}$ where $0\geq j\leq 2n$:
Figure 4.4: $A_{n}$ with orientation and partial labels
Proposition 4.5
The coloring space of $Q(*)$ on $A_{n}$ is the solution set $\{x_{i}=c_{1},y_{1}=y_{2}=c_{2},others=c_{3}~{}|~{}c_{1},c_{2},c_{3}\in Q,~{}c_{1}\sim c_{2}\sim c_{3}\}$. $others$ refer to all $m_{i}$ and all arcs in each $T_{j}$.
Proof for proposition 4.5
$\forall i\in\{1,3,...,2n-1\}$, the pair of crossings $c_{i}$ and $c_{i+1}$ imply $m_{i-1}=m_{i}*^{-1}x_{i}=m_{i+1}$. Therefore we have $m_{0}=m_{2},m_{2}=m_{4},...$, which gives $m_{0}=m_{2}=m_{4}=...=m_{2n}$ ($I$). With this result, the pair of crossings $d_{i}$ and $d_{i+1}$ give $x_{i-1}=x_{i}*m_{i-1}=x_{i}*m_{i+1}=x_{i+1}$, so similarly we have $x_{0}=x_{2}=x_{4}=...=x_{2n}$ ($II$), particularly, $x_{0}=x_{2n}$. For each $j\in\{2,4,...,2n\}$,the crossing $d_{j}$ gives $x_{j-1}=x_{j}*^{-1}m_{j}$. Combining ($I$) and ($II$), we have $x_{j}*^{-1}m_{j}$ is the same for all $j$,so $x_{j+1}$ is also all the same, namely $x_{1}=x_{3}=...=x_{2n-1}$ ($III$)
By using the fact that relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*b=a$ is an equivalence relation, we can then repeat the process in section 4.4 of calculating the crossing equations given by ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}t_{5},t_{6},t_{7},t_{8}}$ in figure 4.4 to derived that $x_{0}=x_{2n}=x_{2n+1}$, $y_{1}=y_{2}$(1), and $m_{0}=m_{2n}$, and $x_{0}\sim y_{1}\sim m_{0}$. Combining this with ($I$), we have $m_{i}\sim y_{1}\sim x_{0}$(2) for all even $i$.
By (2), $x_{0}\sim m_{0}$ so $x_{0}*m_{0}=x_{0}$, so crossing $d_{1}$ gives $x_{1}=x_{0}*^{-1}m_{0}=x_{0}$. Combining with $(II)$ and $(III)$, we have $x_{i}=x_{j}~{}\forall i,j\in\{1,2,...,2n+1\}$(3). For each odd $j$, $j-1$ is even, so by (2) and (3) $m_{j-1}\sim x_{j}$, so the crossing $c_{j}$ gives $m_{j}=m_{j-1}*x_{j}=m_{j-1}$. This means $m_{0}=m_{1},m_{2}=m_{4},...,m_{2n-2}=m_{2n-1}$, and by ($I$), we also have $m_{2n-1}=m_{2n}$.
When the system of crossing relations given by all $T_{i}$ are added with the equations $\{m_{0}=m_{1},m_{2}=m_{4},...,m_{2n-1}=m_{2n}\}$ we just derived, this given the system of crossing relations generated by the knot formed by connecting all $m_{i}$ and $m_{i+1}$ together for $i\in\{0,2,...,2n\}$. This knot is obviously equivalent to an unknot, so the middle complex component in $A_{k}$ degenerate to an unknot under the coloring by $Q(*)$, which means all $m_{i}$ and all arcs in $T_{i}$ must be colored to be the same element, so we can say they also equal to $m_{0}$. This result along with (1), (2) and (3) complete the proof of proposition 4.5.
4.6 summarization
By section 4.1,4.4 and 4.5, if $Q(*)$ is a distributive quandle such that the relation $\sim$ on $Q$ defined by $a\sim b\Leftrightarrow a*c=a$ is an equivalence equation, the solutions sets correspond to the coloring space for both connected sum of two Hopf links $L_{2H}$ and general Allen-Swenberg links $A_{k}$ are in the form $\{x=c_{1},y=c_{2},z=c_{3}~{}\forall x\in X,y\in Y,z\in Z~{}|~{}c_{1}\sim c_{2}\sim c_{3}\in Q(*)\}$, where $X,Y,Z$ is a partition of the correspond fundamental quandle of the link $Q(L)$. This means the coloring space both $L_{2H}$ and any $A_{k}$ should be $\{f:Q(L)\to Q(*)~{}|~{}f(X)=\{c_{1}\},f(Y)=\{c_{2}\},f(Z)=\{c_{3}\},c_{1}\sim c_{2}\sim c_{3}\}$.
Let $X,Y,Z$ be the partition for $L_{2H}$ and $X^{\prime},Y^{\prime},Z^{\prime}$ be the partition for $A_{k}$ respectively in their solution sets defined above. I would first show that there exist a bijection function $f:Hom(L_{2H},Q(*))\to Hom(A_{k},Q(*))$. Let $h\in Hom(L_{2H},Q(*))$ be arbitrary. By the discussion above, $h$ is defined by $h(X)=\{c_{1}\},h(Y)=\{c_{2}\},f(Z)=\{c_{3}\}$, for some $c_{1},c_{2},c_{3}\in Q(*)$ such that $c_{1}\sim c_{2}\sim c_{3}$. So we can define $f$ by defining $f(h):A_{k}\to Q$ to be:
$f(h)(X^{\prime})=\{c_{1}\},f(h)(Y^{\prime})=\{c_{2}\},f(h)(Z^{\prime})=\{c_{3}\}$. $f$ is obviously bijective, which shows $C_{A_{k}}=C_{2H}$. Since $Im(h)=\{c_{1},c_{2},c_{3}\}=Im(f(h))$, this function also preserve the cardinality of the image of the input function.
To show that $\Phi(A_{i},Q(*))=\Phi(L_{2H},Q(*))$ is equivalent of showing that $\forall i\in\{1,2,3\}$,
|$|\{h\in Hom(L_{2H},Q(*)):~{}|Im(h)|=i\}|=|\{g\in Hom(A_{k},Q(*)):~{}|Im(g)|=i\}|$. It suffices to show there is a bijective map between this two set. Obviously, the function $f$ we previously defined restricted to the given domain, $f_{|\{h\in Hom(L_{2H},Q(*))~{}|~{}|Im(h)|=i\}}$, is bijective and indeed a function between this two sets. Therefore, $\Phi(A_{i},Q(*))=\Phi(L_{2H},Q(*))$.
This completes the proof for Proposition 3.1.
5 Reference
[1]V. Chernov and S. Nemirovski. Legendrian links, causality, and the low conjecture. Geom. Fuct. Anal., 19(0222503):1323–1333, 2010.
[2]R. J. Low. Causal relations and spaces of null geodesics. PhD thesis, Oxford University, 1988.
[3]S.Allen and J.Swenberg. Do link polynomials detect causality in globally hyperbolic space-times? J. Math. Phys., 62(3), 2021.
[4] Miller, Jacob. ”Quandle Invariants of Knots and Links.” (2022).
[5] Nelson, Sam. ”Classification of finite Alexander quandles.” arXiv preprint math/0202281 (2002).
[6] A. Navas and S. Nelson. On symplectic quandles. Osaka J. Math, 45(4):973–985, December 2008.